OpenStudy (anonymous):

Solve for x: |x + 2| + 16 = 14. a).x=-32 and x= -4 b).x= -4 and x=0 c).x=0 and x=28 d). no solution

OpenStudy (texaschic101):

| x + 2 | + 16 = 14 we want to get the absolute value sign (with everything in it) on one side of the equal sign, and everything else on the other side. So we would subtract 16 from both sides. | x + 2 | = 14 - 16 | x + 2 | = -2 Now we ser it up like this.. x + 2 = -2 and -(x + 2) = -2.....solving for x giving us 2 answers. x + 2 = -2 x = -2 - 2 x = -4 and -(x + 2) = -2 --- distribute -x - 2 = -2 -x = -2 + 2 -x = 0 x = 0 so your 2 solutions are : x = -4 and x = 0